A 423.1-N force accelerates a go-cart and its driver from 11.3 m/s to 18.3 m/s in 5.01 s. What is the mass of the go-cart plus driver?
1
Expert's answer
2012-09-25T11:06:09-0400
According to the second Newton's law

F = Ma,

where M is the mass of the go-cart with driver. We know that acceleration a is

a = (18.3[m/s] - 11.3[m/s])/5.01[s] ≈ 1.4 m/s².

So, the mass of the go-cart plus driver is

M = F/a = 423.1[N]/1.4[m/s²] = 302.2 kg.